OpenStudy (anonymous):

I'm doing Rational Equations and i'm struggling can someone please help? Slove m+1/2m+1 = 3/5

OpenStudy (anonymous):

I got -4/15 but that isn't the answer.

OpenStudy (anonymous):

Hmmm... That's what I got too

OpenStudy (anonymous):

Solve k+1/2k-1 = 4/5.

OpenStudy (anonymous):

would someone please help?

OpenStudy (anonymous):

I got 6/5

OpenStudy (anonymous):

Solve 7/w+4 = 5/w+4 - 6/w.